it re 19 A ball is B) same. distance the of a of sms in 5 s you cover at an speed 15 m Amyou B) 2m. walk 21 An archer shoots an reaction to this force is the and bowstring. A) combined weight of the arrow arrows push against the bowstring. feet. of the ground against the archers D) grip of the archers hand on the bow. 22) A push on a 1-kg brick accelerates Neglecting friction accelerating a 10-kg brick it. equally A) one-tenth the amount of force. B) 10 times as much force. 100 times as much force. D just as much force. A) acceleration of the ball. D) none of the above O air resistance acting against the ball. 24) 24) A rock weighs 30 N on Earth and another rock weighs 30 Non the Moon. Which rock has the greater mass? A) the one on Earth B) the one on the Moon D) not enough information 25) You cannot exert a force on a wall A) unless you put your mind to it. B) if the wall resists. unless the wall simultaneously exerts the same amount of force on you 26) The distance a freely falling object falls from rest in one-half second is A) 4 m. D) none of the above 27 Two tennis balls fall through the air from a tall building. One of them is filled with1ead pellets. The 27 ball to reach the ground first is the A) regular ball. same for both B-3
Answer #1

20.correct option is (a), 4 m/s × 5s = 20 m

21. Correct option is b, its reaction force will be arrow's push.

22. Correct option is a, from newtons law, F = ma,

For 1 kg, a = F/1 and for 10 kg , a = F/10

23. Correct option is C, its reaction force will be air resistance

24. Correct option is b, as acceleration due to gravity at moon is one sixth of that on Earth.

25. Correct option is c

26.d, [ -g= d/t2] g= 9.81 ,t =0.5 s

27.c, same for both